The Ruelle-Sullivan map for actions of ℝ^{n}

Johannes Kellendonk^{1} and Ian F. Putnam^{2,}^{∗}

1 Institute Girard Desargues, Université Claude Bernard Lyon 1,

F-69622 Villeurbanne; e-mail: kellendonk@igd.univ-lyon1.fr

2 Department of Mathematics and Statistics, University of Victoria,

Victoria, B. C. V8W 3P4, Canada; e-mail: putnam@math.uvic.ca

February 10, 2005

Abstract

The Ruelle Sullivan map for an ℝ^{n}-action on a compact metric space with invariant probability measure is a graded homomorphism between the integer Cech cohomology of the space and the exterior algebra of the dual of ℝ^{n}. We investigate flows on tori to illuminate that it detects geometrical structure of the system. For actions arising from Delone sets of finite local complexity, the existence of canonical transversals and a formulation in terms of pattern equivariant functions lead to the result that the Ruelle Sullivan map is even a ring homomorphism provided the measure is ergodic.

1 Introduction

We consider a variety of cohomology groups for a continuous ℝ^{n}-action φ on a compact metric space X. Among them are the Cech cohomology Ȟ(X, ℤ) of X and the dynamical cohomology of the dynamical system (X, φ) by which we mean the Lie-algebra cohomology H(ℝ^{n}, C^{∞}(X, ℝ)) of ℝ^{n} with coefficients in the φ-smooth real functions on X. An analog of the Cech-de Rham complex provides us with a graded ring homomorphism between the two.

∗Supported by the Natural Sciences and Engineering Research Council Canada

Cohomology captures topological information about the dynamical system. If we are also given an invariant Borel probability measure μ then we can capture some of the geometric structure which comes from the (euclidean) geometry of the ℝ^{n}-orbits. We do this by exploiting Ruelle-Sullivan currents associated with sub-actions. These yield linear functionals on dynamical cohomology which we combine into a graded homomorphism from the cohomology to the exterior algebra of the dual ℝ^{n}^{∗} of ℝ^{n}. The Ruelle Sullivan map

τ_{φ,μ} : Ȟ(X, ℤ) → Λℝ^{n}^{∗}  (1)

is defined as the composition of the two homomorphisms. A similar group homomorphism related to the above but between the K-group K(C(X)) and Λℝ^{n}^{∗} can be found in [Co80].

Our aim here is to study some properties of the Ruelle-Sullivan map and to provide interesting and computable examples which illuminate that the Ruelle-Sullivan map detects geometrical structure. For instance, it captures the continuous eigenvalues of the action (Theorem 14) and can be used to distinguish cut & projection sets of fixed dimension and codimension, which have the same cohomology group but project out of geometrically different lattices (Theorem 15).

A large part of the paper is devoted to the more special class of ℝ^{n}- actions arising from Delone sets of finite local complexity in ℝ^{n}. In this case X = Ω_{P} is the continuous hull of such a Delone set P. This system has the particular advantage that P also defines canonical local transversals for the ℝ^{n}-action. These transversals are used to define H(ℝ^{n}, C_{tlc}^{∞}(Ω_{P} , ℝ)), the transversally locally constant dynamical cohomology of Ω_{P}. It arises from a subcomplex of the Lie-algebra complex and is isomorphic to the Cech cohomology Ȟ(Ω_{P} , ℝ) with real coefficients under the above mentionned ring homomorphism.

Two further groups, Ȟ_{P} (ℝ^{n}) and Ȟ(ℝ^{n}), the strongly and weakly P - equivariant cohomology groups of ℝ^{n}, are defined as cohomologies of subcomplexes of the de Rham complex for ℝ^{n}. Ȟ_{P} (ℝ^{n}) is a more intuitive picture of H(ℝ^{n}, C_{tlc}^{∞}(Ω_{P} , ℝ)). Theorems 20,23 contain the proof of the result announced in [Ke03] that the (strongly) P -equivariant cohomology is isomorphic to the Cech cohomology (with real coefficients) of the continuous hull of P. H_{P} (ℝ^{n}) is the closure of Ȟ_{P} (ℝ^{n}) in an appropriate sense and is isomorphic to the dynamical cohomology H(ℝ^{n}, C^{∞}(Ω_{P} , ℝ)). We use techniques from differential topology to show that the Ruelle-Sullivan map is even a ring homomorphism for ergodic ℝ^{n}-actions associated with Delone sets of finite local complexity (Theorem 26).

2 Ruelle-Sullivan currents

We consider (X, φ), a topological ℝ^{n}-dynamical system. That is, X is a compact metric space and φ is a continuous action of ℝ^{n} on X. For each v in ℝ^{n}, φ_{v} : X → X is a homeomorphism of X and the map sending (x, v) to v ↦ φ_{v}(x) is jointly continuous. Moreover, we have φ_{v} ∘ φ_{w} = φ_{v+w}, for all v, w in ℝ^{n}. We also call (X, φ) simply an ℝ^{n}-action.

We also let μ be a φ-invariant Borel probability measure on X. Such measures exist always [Gl76]. In some cases we will assume that the system is ergodic; that is, a closed invariant set has either measure 1 or measure 0.

We let C(X) denote the algebra of continuous complex functions on X. We call f ∈ C(X) continuously φ-differentiable if

v · f = lim_{t→0} \frac{f(φ_{tv}(x)) − f(x)}{t},  (2)

exists and lies in C(X) for all x ∈ X and v ∈ ℝ^{n}. We say that f is φ-smooth or just smooth if it is infinitely continuously φ-differentiable. We let C^{∞}(X) denote the set of smooth functions in C(X). An easy adaptation of the usual arguments involving convolution with a smooth bump function shows that C^{∞}(X) is uniformly dense in C(X).

More generally, if W is any finite dimensional real vector space, we let C(X, W) denote the set of continuous W -valued functions on X. The definition of C^{∞}(X, W) extends easily.

We let ℝ^{n}^{∗} denote the dual space of the vector space ℝ^{n} and for v ∈ ℝ^{n} and w ∈ ℝ^{n}^{∗}, we denote the pairing by ⟨v, w⟩. We let Λℝ^{n}^{∗} denote the graded exterior algebra of ℝ^{n}^{∗}. We consider C^{∞}(X, Λℝ^{n}^{∗}), which is also a graded ring under point-wise wedge-product, and define d : C^{∞}(X, ℝ) → C^{∞}(X, ℝ^{n}^{∗}),

df(x)(v) = v · f

v ∈ ℝ^{n}. This extends as usual to a differential

d : C^{∞}(X, Λ^{k}ℝ^{n}^{∗}) → C^{∞}(X, Λ^{k+1}ℝ^{n}^{∗}),

i.e. a derivation of degree one with d^{2} = 0. For concreteness, if {e_{1}, …, e_{n}} is a basis for ℝ^{n}, we let {de_{1}, …, de_{n}} denote the dual basis. Then every element of degree k may be written in the form

∑_{I} ω_{I}de_{i_{1}}de_{i_{2}} ⋯ de_{i_{k}},

where I = {i_{1}, …, i_{k}} ⊂ {1, 2, …, n} and ω_{I} ∈ C^{∞}(X, ℝ). Furthermore,

dω = ∑_{i=1}^{n} \frac{∂ω}{∂e_{i}} de_{i},  \frac{∂ω}{∂e_{i}} = lim_{t→0} \frac{ω(φ_{te_{i}}(x)) − ω(x)}{t} .  (3)

Definition 1. The Lie-algebra cohomology of ℝ^{n} with values in C^{∞}(X, ℝ), H(ℝ^{n}, C^{∞}(X, ℝ)), is the cohomology of the complex (C^{∞}(X, Λℝ^{n}^{∗}), d). We call is also simply the dynamical cohomology of (X, φ).

The product on C^{∞}(X, Λℝ^{n}^{∗}) induces a graded ring structure on H(ℝ^{n}, C^{∞}(X, ℝ)). If φ is locally free, i.e. if for all x ∈ X there is ϵ such that v < ϵ and φ_{v}(x) = x imply v = 0, then X is a foliated space in the sense of [MS88] and the above definition coincides with their definition of tangential cohomology. However we do not require the local freeness of φ.

We may use our invariant measure μ to define a map from this cohomology to Λℝ^{n}^{∗}.

Definition 2. The Ruelle-Sullivan current C_{μ} associated with μ is the linear map

⟨C_{μ}, ·⟩ : C^{∞}(X, Λℝ^{n}^{∗}) → Λ^{k}ℝ^{n}^{∗},

defined by

⟨C_{μ}, ω⟩ = ∫_{X} ω(x)dμ(x).

Lemma 3. Let μ be an invariant probability measure for the action φ. Let f be any φ-smooth W -valued function. Then

⟨C_{μ}, dω⟩ = 0,

for all ω in C^{∞}(X, Λℝ^{n}^{∗}).

Proof. This follows from (3) together with the invariance of μ under φ.

Corollary 4. The pairing with C_{μ} descends to a map on cohomology,

τ̃_{φ,μ} : H(ℝ^{n}^{∗}, C^{∞}(X, ℝ)) → Λℝ^{n}^{∗}.

Remark 5. All of this section can be easily generalized to the case of an action φ of an arbitrary Lie-group G on X. Since this would be the right framework for tilings which have finite local complexity w.r.t. a larger subgroup of the Euclidean group than the translation group we indicate the main changes. The definition of φ-differentiable (2) is expressed with the help of the Lie-algebra action involving the exponential map exp : G → LieG, namely v·f := lim_{t→0} \frac{f(φ_{exp tv}(x))−f(x)}{t}. The differential d : C^{∞}(X, Λ^{k}LieG^{∗}) → C^{∞}(X, Λ^{k+1}LieG^{∗}) is defined by

df(v) = v · f ,
dw(v_{1}, v_{2}) = −w([v_{1}, v_{2}]) ,

for f ∈ C^{∞}(X, Λ^{0}LieG^{∗}) and w ∈ LieG^{∗} viewed as constant smooth function in C^{∞}(X, Λ^{1}LieG^{∗}), v, v_{1}, v_{2} ∈ LieG. Finally, the Ruelle-Sullivan current takes values in ΛLieG^{∗} and satisfies ⟨C_{μ}, dω⟩ = d⟨C_{μ}, ω⟩. Hence it descends to a map on the Lie-algebra cohomology:

τ̃_{φ,μ} : H(LieG, C^{∞}(X)) → H(LieG, ℝ).

3 The Cech-cohomology and the Ruelle-Sullivan map

The purpose of this section is to construct a morphism from the Cech cohomology Ȟ(X, ℤ) of X to the dynamical cohomology of (X, φ) and to prolong τ̃_{μ,φ} to a map τ_{μ,φ} : Ȟ(X, ℤ) → Λℝ^{n}^{∗}, the Ruelle-Sullivan map.

We begin by fixing a open cover of X, U = {U_{i}}_{i}, and a partition of unity ρ_{i} which is subordinate to this cover, i.e. suppρ_{i} ⊂ U_{i}. For i_{0} < i_{1} < … < i_{j}, we let

U_{i_{0}i_{1}···i_{j}} = U_{i_{0}} ∩ U_{i_{1}} ∩ ⋯ U_{i_{j}}.

We note that for any open set U, any continuous function f on U and vector v in ℝ^{n}, the expression v · f still makes sense as a function on U (assuming the limit exists) even though the set U is not invariant under φ. So we define C^{∞}(U, Λℝ^{n}^{∗}) as the smooth functions on U taking values in Λℝ^{n}^{∗}.

We define the double complex

K^{j,k}(U) = ⨁_{i_{0}<i_{1}<···<i_{j}} C^{∞}(U_{i_{0}i_{1}···i_{j}}, Λ^{k}),

where here and below the sum is only over i_{0} < i_{1} < ⋯ < i_{j} with non-empty U_{i_{0}i_{1}···i_{j}}. The i_{0} ⋯ i_{j}-component of ω ∈ K^{j,k}(U) is denoted by ω_{i_{0}···i_{j}}. For notational purposes, we also define ω_{i_{0}···i_{j}} = 0 if U_{i_{0}i_{1}···i_{j}} = ∅, and ω_{σ(i_{0})···σ(i_{j})} = sgn(σ)ω_{i_{0}···i_{j}}, for any permutation σ.

The double complex has two commuting differentials, the Lie-algebra differential d : K^{j,k}(U) → K^{j,k+1}(U), defined above and the Cech differential

δ : K^{j,k}(U) → K^{j+1,k}(U),

given by

(δω)_{i_{0}···i_{j+1}} = ∑_{l=0}^{j+1}(−1)^{l}ω_{i_{0}···î_{l}···i_{j+1}},

î_{l} meaning that i_{l} is omitted.

Using the subordinate partition of unity {ρ_{i}}_{i} we define for j > 0, h : K^{j,k}(U) → K^{j−1,k}(U),

(hω)_{i_{0}···i_{j−1}} = ∑_{i} ρ_{i}ω_{ii_{0}···i_{j−1}}.

Then hδ + δh = 1 and so for any k the complex (Mayer-Vietoris sequence)

0 → C^{∞}(X, Λ^{k}ℝ^{n}^{∗}) →^{r} K^{0,k}(U) →^{δ} ⋯ K^{j,k}(U) →^{δ} K^{j+1,k}(U) ⋯ (4)

is exact, i.e. has trivial cohomology, where r(ω)_{i} is the restriction of ω to U_{i}.

We let Č(U_{i_{0}i_{1}···i_{j}}, ℝ) and Č(U_{i_{0}i_{1}···i_{j}}, ℤ) denote the functions from U_{i_{0}i_{1}···i_{j}} to ℝ and ℤ, respectively, which are locally constant, and hence smooth. Then we set

Č^{j}(U, ℝ)) = ⨁_{i_{0}<i_{1}<···<i_{j}} Č(U_{i_{0}i_{1}···i_{j}}, ℝ),
Č^{j}(U, ℤ)) = ⨁_{i_{0}<i_{1}<···<i_{j}} Č(U_{i_{0}i_{1}···i_{j}}, ℤ).

It is clear that we have

Č^{j}(U, ℤ) ⊂ Č^{j}(U, ℝ) ⊂ K^{j,0}(U),

and together with the map δ, these form subcomplexes. We let i denote the inclusion of either of the first two in the last. The cohomology (in δ) of the subcomplexes no longer vanishes, because the corresponding maps h involve functions (the ρ_{i}) which are not locally constant. Then, Ȟ(U, ℝ) and Ȟ(U, ℤ) are the Cech cohomologies of the covering U with coefficients in ℝ and ℤ, respectively. For notational convenience, we will use Č^{j}(U) and Ȟ(U) to denote either of Č^{j}(U, ℝ) or Č^{j}(U, ℤ) and the corresponding cohomology.

Ȟ(U) carries a graded ring structure [BT82]. The product is induced from the map Č^{j}(U) × Č^{k}(U) → Č^{j+k}(U), (ω, ω′) ↦ ω · ω′,

(ω · ω′)_{i_{0}···i_{j+k}} = (−1)^{jk}ω_{i_{0}···i_{j}} ω′_{i_{j+1}···i_{j+k}}.

It is useful to have the following special case in mind. If φ is a free ℝ^{n}-action and X consists of a single orbit then the action induces on X the structure of a differentiable manifold and the Lie-algebra cohomology with coefficients in C^{∞}(X, ℝ) agrees with its de Rham cohomology. The above double-complex is then the Cech-de Rham double complex used to construct a ring homomorphism between the Cech cohomology of X and its de Rham cohomology [BT82]. When Cech cohomology with real coefficients is considered this homomorphism becomes an isomorphism.

In our more general situation the same approach yields a ring homomorphism between the Cech cohomology of X and the dynamical cohomology of (X, φ), but it won’t, in general, become an isomorphism when real coefficients are considered. This is the next result, whose proof goes exactly as in the de Rham theorem [BT82]. Note that h maps K^{j,k}(U) to K^{j−1,k}(U), while d maps K^{j−1,k}(U) to k^{j−1,k+1}(U). Hence, (dh)^{n} will map K^{n,0}(U) to K^{0,n}(U), for all n ≥ 0.

Theorem 6. The maps

(−1)^{j}r^{−1}(dh)^{j}i : Č^{j}(U) → C^{∞}(U, Λ^{j}),

induce a graded ring homomorphism

θ_{φ,U} : Ȟ(U) → H(ℝ^{n}, C^{∞}(X, ℝ)),

both, with coefficients ℤ or ℝ for the Cech cohomology.

Recall that an open cover U′ is a refinement of U, if there is a map α : U′ → U such that U ⊂ α(U). Such a map induces a map at the level of cohomology Ȟ(U) → Ȟ(U′) which is independent of the choice of α. The Cech cohomology of X, Ȟ(X) is defined to be the inductive limit of the groups Ȟ(U) over all open covers U of X. The morphisms θ_{φ,U} furnish a graded ring homomorphism θ_{φ} : Ȟ(X) → H(ℝ^{n}, C^{∞}(X, ℝ)), both, with coefficients ℤ or ℝ for the Cech cohomology.

Definition 7. Let (X, φ) be an ℝ^{n}-action with φ-invariant measure μ. The Ruelle-Sullivan map τ_{φ,μ} : Ȟ(X, ℤ) → Λℝ^{n}^{∗} is defined by

τ_{φ,μ}(a) = ⟨C_{μ}, θ_{φ}(a)⟩.

In particular, if a is represented as an element of Ȟ(U, ℤ), where U is an open cover of X, then τ_{φ,μ}(a) = ⟨C_{μ}, θ_{φ,U}(a)⟩.

The Ruelle-Sullivan map does not have to be injective. But it contains information coming from an invariant measure and therefore about the (euclidean) geometry of the orbits. The philosophy is that Cech cohomology together with the Ruelle-Sullivan map will furnish a better invariant for ℝ^{n}- actions.

Clearly, the Ruelle-Sullivan map is a graded homomorphism of groups. We will discuss in Section 6.1 a large class of examples for which it is even a graded ring homomorphism.

Remark 8. Let us comment on the terminology. This is really just for motivation, so we will not go into any details. Given a foliation and a transverse invariant measure, Ruelle and Sullivan constructed a current, usually called the Ruelle-Sullivan current. In our situation we have an action rather than just a foliation but if the action is locally free this yields a foliated space, in the sense of Moore and Schochet [MS88]. However, we have much more. If one selects a linearly independent set of k vectors, one can define an action of ℝ^{k} by restricting the flow to the linear span of the vectors. Any invariant measure for the ℝ^{n}-action can be used to give an invariant measure for this action. Associated to this is a Ruelle-Sullivan current which yields a map from cohomology of X to the real numbers. If one fixes the cohomology class and considers the set of vectors to vary, this is an element of degree k in the exterior algebra of ℝ^{n}^{∗}. Hence, the Ruelle-Sullivan construction can be viewed as giving map from the cohomology to the exterior algebra of ℝ^{n}^{∗}. This is our Ruelle-Sullivan map.

Remark 9. Also the double complex generalises to the case of an arbitrary Lie-group action on X. As a result one gets a Ruelle-Sullivan map

τ_{φ,μ} : Ȟ(X, ℤ) → H(LieG, ℝ).

Remark 10. Connes presents in [Co80] a similar group homomorphism in non-commutative geometry. He starts with a Lie-group G action on a C^{∗}- algebra A with invariant trace tr to construct a group homomorphism from the K-group of A to the cohomology of the Lie-algebra,

Ch_{tr} : K_{i}(A) → ⨁_{k} H^{i+2k}(LieG, ℝ).

When A = C(X) with trace induced from an invariant measure μ and rational coefficients are considered, Connes’ construction specialises to the one considered above; Ch_{tr} factors through the Chern character identifying K(C(X)) ⊗ ℚ with Ȟ(X, ℚ). The question when this identification can be made even over integer coefficients is a lot harder.

Coming back to G = ℝ^{n} where the Connes-Thom isomorphism yields K_{i}(C(X)) ≅ K_{i−n}(C(X) ⋊_{φ} ℝ^{n}) the components of the Ruelle Sullivan map may be also be related to functionals on K(C(X) ⋊_{φ} ℝ^{n}) which arise from pairings with certain cyclic cocycles of C^{∞}(X) ⋊_{φ} ℝ^{n} [C94]. We do not elaborate on that apart from mentionning that for a minimal, uniquely ergodic dynamical system (X, φ) whose action is locally free and such that X has a totally disconnected transversal we have, at least if Ȟ(X, ℤ) is torsion free, that K_{0}(C(X) ⋊_{φ} ℝ^{n}) ≅ ⨁_{k} Ȟ^{n+2k}(X, ℤ) [FH99] and the state on K_{0}(C(X) ⋊_{φ} ℝ^{n}) defined by the unique measure on X coincides when restricted to Ȟ^{n}(X, ℤ) with the restriction of τ_{φ,μ} to Ȟ^{n}(X, ℤ). For these systems the Ruelle-Sullivan map is therefore an extension of the gap-labelling.

We end this section with two results which will be useful for our computations.

Proposition 11. Let (X, ψ) and (Y, φ) be ℝ^{n}-actions with invariant measures μ and ν, respectively. Suppose that

η : Y → X

is a continuous function which is measure preserving and equivariant; i.e. μ = ν ∘ η^{−1} and φ_{v} ∘ η = η ∘ ψ_{v}, for all v in ℝ^{n}. Then

Ȟ(X, ℤ) ⟶^{τ_{ψ,ν}} Λℝ^{n}^{∗}
η^{∗}↓  ↓^{=}
Ȟ(Y, ℤ) ⟶^{τ_{φ,μ}} Λℝ^{n}^{∗}

is a commutative diagram.

Proof. A straightforward substitution in the integral.

We also want to understand how our map is affected by linear reparametrizations of the flow.

Proposition 12. Let (X, φ) be an ℝ^{M}-action and A : ℝ^{n} → ℝ^{M} be a linear map. Define an ℝ^{n} action, ψ, on X by ψ_{v} = φ_{Av}. Then

Ȟ(X, ℤ) ⟶^{τ_{φ,μ}} Λℝ^{M}^{∗}
=↓  ↓^{ΛA^{∗}}
Ȟ(X, ℤ) ⟶^{τ_{ψ,μ}} Λℝ^{n}^{∗}

is a commutative diagram where ΛA^{∗} denotes the extension of the dual map A^{∗} : ℝ^{M}^{∗} → ℝ^{n}^{∗} to the exterior algebras.

Proof. For clarity we denote the differentials by d_{φ} and d_{ψ}. Let v ∈ ℝ^{M}. Then by (3) ⟨v, A^{∗}d_{φ}f⟩ = ⟨Av, d_{φ}f⟩ = ⟨v, d_{ψ}f⟩. Hence A^{∗}d_{φ}f = d_{ψ}f which implies θ_{ψ}^{j} = Λ^{j}A^{∗}θ_{φ}^{j} and therefore τ_{ψ,μ} = ΛA^{∗}τ_{φ,μ}.

4 Flows on tori

We discuss a variety of examples which are all related to flows on tori.

Let Γ denote a lattice in ℝ^{M}; that is, Γ is a cocompact discrete subgroup of ℝ^{M}. We consider the torus X_{Γ} = ℝ^{M}/Γ with q : ℝ^{M} → X_{Γ} the quotient map. Let A : ℝ^{n} → ℝ^{M} be a linear map. We define

φ_{v}^{A}(q(x)) = φ_{Av}(q(x)), φ_{v}(q(x)) = q(x + v)  (5)

for all q(x) in X_{Γ} and v in ℝ^{n}. This is our basic dynamical system. We call it the flow on the torus defined by (A, Γ). If it has one dense orbit then it is minimal and uniquely ergodic, the measure being the normalised Lebesgue measure. If the system is not minimal there are other invariant measures. We note that φ^{A} is locally free whenever A is injective. In that case we may call it a generalized Kronecker flow.

Of course, for all choices of A and Γ, Ȟ(X_{Γ}, ℤ) is isomorphic as a graded ring to the exterior ring Λℤ^{M}. What we are after here is the image of the Ruelle-Sullivan map, which will, as it contains geometrical data, distinguish between different choices for Γ and A.

Theorem 13. Consider the flow on the torus defined by (A, Γ) as above. Let μ be the normalised Lebesque measure on X_{Γ}. The corresponding Ruelle-Sullivan map is a surjective graded ring homomorphism from Ȟ(X_{Γ}, ℤ) onto Λ(A^{∗}Γ^{∗}). Here A^{∗} is the dual of A and Γ^{∗} the dual (or reciprocal) lattice. If the restriction of A^{∗} to Γ^{∗} is injective then the Ruelle-Sullivan map is also injective.

Proof. We begin with the case that n = M, A = id, the identity, and Γ = ℤ^{M} the standard cubic lattice generated by an orthormal basis {e_{i}}_{i}. Then the generators of θ_{φ}(Ȟ(X_{ℤ^{M}}, ℤ)) have representatives given by forms of the type dx_{i_{1}} ⋯ dx_{i_{k}} and we see immediately that the latter has image under the Ruelle-Sullivan map de_{i_{1}} ⋯ de_{i_{k}}. Hence the Ruelle-Sullivan map is a graded ring isomorphism between Ȟ(X_{ℤ^{M}}, ℤ) and Λℤ^{M}^{∗}.

Next we consider the case Γ = ℤ^{M} but general A which yields by Proposition 12 that the Ruelle-Sullivan map is a surjective graded ring homorphism between Ȟ(X_{ℤ^{M}}, ℤ) and ΛA^{∗}ℤ^{M}^{∗}.

Finally we consider the case that Γ = Bℤ^{M} for some invertible matrix B and general A. Hence B : ℝ^{M} → ℝ^{M} induces a map X_{ℤ^{M}} → X_{Γ} intertwining the action φ^{A} with φ^{BA}. From Proposition 12 it follows that the Ruelle-Sullivan map is a surjective graded ring homomorphism between Ȟ(X_{Γ}, ℤ) and Λ(A^{∗}B^{∗}ℤ^{M}^{∗}). B^{∗}ℤ^{M}^{∗} = Γ^{∗} is the dual lattice.

If the restriction of A^{∗} to Γ^{∗} is injective then, by multiplicativity, this also the case for the restriction of ΛA^{∗} to ΛΓ^{∗}. This implies by Proposition 12 the injectivity of τ_{φ,μ}.

4.1 Eigenvalues of minimal actions

We consider the above example but in the extreme case that n ≥ M = 1, Γ = ℤ, and A is surjective. We look at X_{ℤ} as the unit circle in the complex plane. There is a non-zero vector w in ℝ^{n}^{∗} such that Av = ⟨v, w⟩. The ℝ^{n}-action on the complex unit circle is thus given by

φ_{v}(z) = e^{2πi⟨v,w⟩}z,  (6)

for v in ℝ^{n}, and z in X_{ℤ}. Lebesgue measure on the circle is the unique invariant probability measure for this action. By Theorem 13 the Ruelle Sullivan map identifies the generator of Ȟ^{1}(X_{Γ}, ℤ) ≅ ℤ with w.

Now let φ be a minimal action of ℝ^{n} on the compact space X with invariant probability measure μ. An element w ∈ ℝ^{n}^{∗} is called a continuous eigenvalue of the action if there is a continuous non-zero function f : X → ℂ such that

f ∘ φ_{v} = e^{2πi⟨v,w⟩}f,

for all v in ℝ^{n}. Solomyak characterises the continuous eigenvalues of a minimal ℝ^{n}-action on a compact metric space (X, φ) with the help of the metric D as follows [So98]: For δ > 0, x ∈ X let Θ_{δ}(x) = {v ∈ ℝ^{n}|D(φ_{v}(x), x) < δ}. Then w is a continuous eigenvalue if and only if for some x ∈ X

lim_{δ→0} sup_{v∈Θ_{δ}(x)} |e^{2πi⟨v,w⟩} − 1| = 0.

Theorem 14. Let φ be a minimal action of ℝ^{n} on the compact metric space X with invariant measure μ. τ_{φ,μ}Ȟ^{1}(X, ℤ) contains all continuous eigenvalues of φ.

Proof. By rescaling f we may assume it has absolute value 1 at some point. From the eigenvalue condition, the set of points where it has absolute value 1 is φ-invariant, and since the flow is minimal, this must be all of X. The map f provides an ℝ^{n}-equivariant map from X to S^{1}, the latter with action as given in (6). The conclusion follows from Proposition 11 with η = f.

4.2 Cut & project sets: window independent cohomological information

We now consider cut & project sets which also come under the name projection point patterns or model sets. We restrict to the case in which the internal space is a real vector space. We will see that part of their cohomology is determined by a generalised Kronecker flow on a torus.

Consider a decomposition of ℝ^{M} into two orthogonal subspaces E and E^{⊥} and a regular lattice Γ ⊂ ℝ^{M}. Denoting π the orthogonal projection onto E and π^{⊥} its orthogonal complement we suppose that the restriction of π to Γ is injective, i.e. E^{⊥} ∩ Γ = {0}, and that π^{⊥}(Γ) is dense in E^{⊥}. These data (E, E^{⊥}, Γ) constitute a cut & project scheme as in [Mo97]. Identifying the dual of E and E^{⊥} (as abelian groups) with E and E^{⊥}, respectively, the data (E, E^{⊥}, Γ^{∗}) define the dual cut & project scheme. Indeed, the restriction of π to Γ being injective implies denseness of π^{⊥}(Γ^{∗}), and denseness of π^{⊥}(Γ) implies that the restriction of π to Γ^{∗} is injective [Mo97].

To construct a so-called cut & project set in E one is also given subset K ⊂ E^{⊥} which is compact and the closure of its interior. Then

P_{K} = {π(γ)|π^{⊥}(γ) ∈ K}

is called the cut & project set with window K. P_{K} and K are called nonsingular if π^{⊥}(Γ) ∩ ∂K = ∅.

As we will describe in the next section in more detail, P_{K} defines an ℝ^{n}- action on a topological space Ω_{P_{K}}, called the hull of P_{K}. For non-singular P_{K}, the hull has the form Ω_{P_{K}} = (E ⊕ E_{c}^{⊥})/Γ where E_{c}^{⊥} is the cut up internal space. It is obtained from E^{⊥} by disconnecting it along all points of ∂K + π^{⊥}(Λ). Although the latter is a dense set, the resulting surjection η : E_{c}^{⊥} → E^{⊥} is almost one-to-one. The ℝ^{n}-action on Ω_{P_{K}} which is induced from the translation of P_{K} in ℝ^{n} (see Section 5) is here simply given by ψ_{v}(q̃(x, y)) = q̃(x + v, y) where q̃: E ⊕ E_{c}^{⊥} → (E ⊕ E_{c}^{⊥})/Γ is the canonical projection. η induces an almost one-to-one surjection η : Ω_{P_{K}} → X_{Γ} = ℝ^{M}/Γ interwining ψ with the flow associated to (A, Γ) where, in the notation above, ℝ^{n} = E and A is the inclusion map E ⊂ ℝ^{M}. The generalised Kronecker flow is minimal and uniquely ergodic and the uniquely ergodic measure on Ω_{P_{K}} is the pull back of the measure on the Kronecker flow under η [FHK02].

Our aim is to study here the part of the cohomology which does not depend on the cuts but only on the cut & projection scheme. This is, more precisely, the image of η^{∗} : Ȟ(X_{Γ}, ℤ) → Ȟ(Ω_{P_{K}}, ℤ).

Theorem 15. Consider a cut & project set as above. Then the Ruelle-Sullivan map associated with its ℝ^{n}-action restricts to a graded ring isomorphism between η^{∗}(Ȟ(X_{Γ}, ℤ)) and Λπ(Γ^{∗}).

Proof. Application of Proposition 11 reduces the system to the Kronecker flow associated to (A, Γ) where A is the inclusion E ⊂ ℝ^{M}. Now the dual of A is precisely the projection π whose restriction to Γ^{∗} is injective. The statement follows therefore from Theorem 13.

If also the restriction of π^{⊥} to Γ is injective, i.e. E ∩ Γ = {0}, then P_{K} has no periods, and vice versa. This is also equivalent to the denseness of π(Γ^{∗}) [Mo97]. We are thus led to the following conjecture. The definition of a Delone set of finite local complexity will be given in the next section.

Conjecture 16. Let P be a Delone set of finite local complexity. Suppose that its associated ℝ^{n}-action is minimal and uniquely ergodic. Then ℝ^{n}^{∗}/\overline{τ_{φ,μ}Ȟ^{1}(Ω_{P} , ℤ)} is the dual of the lattice of periods of P.

5 Transversally locally constant forms and hulls of aperiodic systems

We now turn to the issue of a de Rham type isomorphism. We mentionned that the dynamical cohomology of (X, φ) coincides with the usual de Rham cohomology in the case that X consists only of one orbit and φ is locally free. In this case the de Rahm theorem applies which states that when real coefficients are considered the homomorphism θ_{φ} of Theorem 6 becomes an isomorphism for good covers. This cannot be expected for the general case. Our main applications are to the study of aperiodic order and more specifically aperiodic tilings or point patterns of Euclidean space. In that case we will describe the appropriate refinement of dynamical cohomology which is needed to obtain an isomorphism.

We find it most convenient to work in the context of Delone sets but, as is well known, we could equally well talk about tilings. Recall that a set P in ℝ^{n} is a Delone set if there are positive constants r_{P} , R_{P} such that

1. P is r_{P} -discrete, i.e. for all x ≠ y in P, |x − y| ≥ r_{P}, and

2. for all x in ℝ, there is y in P with |x − y| ≤ R_{P}.

We say that P is aperiodic if, for any x in ℝ^{n}, P + x = P holds only if x = 0. Finally, P has finite local complexity if, for any fixed R, the number of sets (P − x) ∩ B(0, R), for x in P, is finite. Here B(x, R) is the (open) ball of radius R around x. All Delone sets considered in this article are assumed to have finite local complexity. It is worth stating that the proper notion of finite local complexity is relative to a subgroup of the isometry group of ℝ^{n}. Here, we are implicitly using the subgroup of translations. When more general sub-Lie groups of the isometry group are considered the generalisations indicated in Remarks 5,9 have to be taken into account.

The hull associated with P, denoted by Ω_{P}, is a compact metric space obtained by completion of the set of all translations of P, {P + x | x ∈ ℝ^{n}}, in the metric

D(P, P′) = inf\{ϵ | ∃x, x′ ∈ B(0, \frac{ϵ}{2}) : (P −x)∩B(0, ϵ^{−1}) = (P′−x′)∩B(0, ϵ^{−1})\}.

The elements of Ω_{P} can all be interpreted as Delone sets. The natural translation of Delone sets in ℝ^{n} extends to an action of ℝ^{n} on Ω_{P}. This is our action: φ_{v}(P′) = P′ − v, for P′ in Ω_{P} and v in ℝ^{n}.

A fair amount is now known about the structure of such spaces. One particularly interesting feature is that the space possesses canonical local transversals to the action of ℝ^{n}. For each P′ in Ω_{P} and ϵ > 0, we define

T_{P′,ϵ} = {P′′ ∈ Ω_{P} | P′′ ∩ B(0, ϵ^{−1}) = P′ ∩ B(0, ϵ^{−1})}.

These closed sets are transversal in the sense that P′′ ∈ T_{P′,ϵ} implies φ_{v}(P′′) ∉ T_{P′,ϵ} if 0 ≠ |v| < \frac{r_{P}}{2} .

Definition 17. A function f defined on Ω_{P} is transversally locally constant if, for every P′ in Ω, there is ϵ > 0 such that f is constant on T_{P′,ϵ}. C_{tlc}^{∞}(Ω_{P} , Λ) denotes the transversally locally constant functions in C^{∞}(Ω_{P} , Λ). The Lie-algebra cohomology H(ℝ^{n}, C_{tlc}^{∞}(Ω_{P} , ℝ)) of ℝ^{n} with coefficients in C_{tlc}^{∞}(Ω_{P} , Λ) will also be called the transversally locally constant dynamical cohomology of Ω_{P}.

For any open cover U of Ω_{P} and j, k ≥ 0, we define K_{tlc}^{j,k}(U) to be those elements of K^{j,k}(U) which are transversally locally constant. It is clear from the definitions that K_{tlc}^{∗}(U) forms a subcomplex of the double complex K^{∗}(U).

In our case, there are natural choices for the open covers U. We let V be an open subset of ℝ^{n} and

T_{P′,ϵ,V} = {P′′ + v : P′′ ∈ T_{P′,ϵ}, v ∈ V }.

Given any R > 0 the sets of the form T_{P′,ϵ,B(x,r)} with x ∈ ℝ^{n} and r ≤ R generate the topology of Ω_{P}.

Lemma 18. Č^{∗}(U, ℝ) is contained in K_{tlc}^{∗,0}(U) and coincides with the kernel of d on K_{tlc}^{∗,0}(U).

Proof. Let V = B(x, r) with r < \frac{r_{P}}{2} . Then T_{P′,ϵ,V} is diffeomorphic to T_{P′,ϵ} × V, the diffeomorphism’s inverse being (P′, v) ↦ P′ + v (here the smoothness of the diffeomorphism should be understood w.r.t. φ-differentiability, in the transverse direction only continuity is required). Under this diffeomorphism a transversally locally constant function is constant in the second coordinate and the differential d becomes the identity times the exterior derivative on ℝ^{n}. Since the topology is generated by the above sets the lemma is thus a direct consquence of the definition of transversally locally constant.

Lemma 19. Let U be a covering by sets T_{P′,ϵ,V} with convex V of diameter less than \frac{r_{P}}{4} . Then the complexes

K_{tlc}^{j,0}(U) →^{d} K_{tlc}^{j,1}(U) →^{d} K_{tlc}^{j,2}(U) →^{d} ⋯

are acyclic. That is, their cohomology vanishes except in degree zero.

Proof. As usual, the result follows once we have found a contracting homotopy h̃ : K^{j,k}→K^{j,k−1}, h̃d + dh̃ = 1, for k ≥ 1.

For that we need to make sure that the covering is good in the sense that finite intersections of the type ⋂_{i} T_{P_{i},ϵ_{i},V_{i}} have contractible leaves. Consider pairs (S, U) where S is a discrete subset of an open set U ⊂ ℝ^{n}. Let T_{S,U} = {P′ ∈ Ω_{P} : P′ ∩ U = S} and T_{S,U,V} = T_{S,U} + V. Then

T_{S,U} ∩ T_{S′,U′} = \{ \begin{matrix}T_{S∪S′,U∪U′} & if U ∩ S′ ⊂ S and U′ ∩ S ⊂ S′ \\ ∅ & otherwise \end{matrix}

and T_{S,U} + v = T_{S+v,U+v}. Using these rules one finds

T_{S,U,V} ∩ T_{S′,U′,V′} = ⋃_{v∈V} ⋃_{w∈V′−v} T (w) + v

where

T (w) = \{ \begin{matrix}T_{S∪(S′+w),U∪(U′+w)} & if U ∩ (S′ + w) ⊂ S \\ & and (U′ + w) ∩ S ⊂ (S′ + w), \\ ∅ & otherwise. \end{matrix}

If the diameter of V and V′ is smaller than \frac{r_{P}}{4} then |w| < \frac{r_{P}}{2} so that, by the r_{P} discreteness of P, T (w) ≠ ∅ for at most one w. For such a w we thus have T_{S,U,V} ∩T_{S′,U′,V′} = T_{S∪(S′+w),U∪(U′+w),V ∩(V′−w)}. We see that V ∩(V′ −w) is convex and hence contractible. Finite intersections ⋂_{i} T_{P_{i},ϵ_{i},V_{i}} fall in the above framework and are thus empty or of the form T_{S,U,V} for some S, U, V with contractible V of diameter smaller than \frac{r}{2}.

In particluar, as in the proof of Lemma 18, ⋂_{i} T_{P_{i},ϵ_{i},V_{i}} = T_{S,U,V} is diffeomorphic to the Cartesian product T_{S,U} ×V. Taking therefore in these product coordinates h̃ = 1 × H, where H is the standard contracting homotopy on ℝ^{n} [AM85], we obtain a map which preserves transversally locally constant functions and satisfies h̃d + dh̃ = 1 × (Hd̃+ d̃H) = 1 (here d̃ is the exterior derivative on ℝ^{n}).

The following theorem follows now as in the classical situation of the de Rham theorem from the last two lemmata together with the observation that any cover of Ω has a refinement by covers of the type used in the last lemma.

Theorem 20. The ring homomorphism θ_{φ} of Theorem 6 provides an isomorphism between Ȟ(Ω_{P} , ℝ), the Cech cohomology of Ω_{P} with real coefficients, and H(ℝ^{n}, C_{tlc}^{∞}(Ω_{P} , ℝ)), the transversally locally constant dynamical cohomology.

6 P -equivariant cohomology

The notion introduced in the last section of transversally locally constant forms on the hull Ω_{P} (and hence of the transversally locally constant dynamical cohomology) can be reduced to something simpler and much less technical. This is the aim of the current section.

We begin by introducing the notion of a strongly P -equivariant function on ℝ^{n} [Ke03] where P is as in the last section a Delone set of ℝ^{n} of finite local complexity. Roughly speaking, a function f on ℝ^{n} is strongly P -equivariant if there is some constant R such that, if the patterns in P surrounding two points x and y of radius R are equal (after translating by −x and −y, respectively), then f must take the same values at x and y. More precisely, we have the following.

Definition 21. Let f be a function defined on ℝ^{n}. We say that f is strongly P -equivariant if there is a constant R > 0 such that, if x, y in ℝ^{n} satisfy

(P − x) ∩ B(0, R) = (P − y) ∩ B(0, R),

then f(x) = f(y).

For a finite dimensional vector space W, we let Č_{P}^{∞}(ℝ^{n}, W) denote the smooth strongly P -equivariant functions from ℝ^{n} to W. The cohomology of the sub-complex of strongly P -equivariant forms on ℝ^{n} is called the strongly P -equivariant cohomology of ℝ^{n} and denoted by Ȟ_{P} (ℝ^{n}).

The functions of the closure C_{P}^{∞}(ℝ^{n}, W) of Č_{P}^{∞}(ℝ^{n}, W) in C^{∞}(ℝ^{n}, W) (w.r.t. the standard Fréchet topology) are called weakly P -equivariant. The cohomology of the sub-complex of weakly P -equivariant forms on ℝ^{n} is called the weakly P -equivariant cohomology of ℝ^{n} and denoted by H_{P} (ℝ^{n}). We note that P -equivariant forms are always bounded.

Proposition 22. 1. Let f : Ω_{P} → W be transversally locally constant. The function f_{P} defined on ℝ^{n} by

f_{P} (x) = f(P − x)

is strongly P -equivariant. In addition, if f is φ-smooth, then f_{P} is a smooth function.

2. If g : ℝ^{n} → W is a strongly P -equivariant function, then the function g̃(P − x) = g(x) extends to a continuous transversally locally constant function on Ω_{P}. If g is smooth, then g̃ is φ-smooth.

3. The map f ↦ f_{P} is a graded ring isomorphism between C_{tlc}^{∞}(Ω_{P} , Λℝ^{n}^{∗}) and Č_{P}^{∞}(ℝ^{n}, Λℝ^{n}^{∗}) intertwining the differential d on the former with the usual exterior derivative on forms on ℝ^{n}. It extends by continuity to a graded ring isomorphism between C^{∞}(Ω_{P} , Λℝ^{n}^{∗}) and C_{P}^{∞}(ℝ^{n}, Λℝ^{n}^{∗}).

Proof. 1. By a partition of unity arguement it suffices to show the statement for a function f which has support in some T_{P′,ϵ,B(0, \frac{r_{P}}{2} )} and on which it is transversally constant. In other words, under identification T_{P′,ϵ,B(0, \frac{r_{P}}{2} )} ≅ T_{P′,ϵ} × B(0, \frac{r_{P}}{2} ), f is of the form f = 1 × g where g : ℝ^{n} → W has support in B(0, \frac{r_{P}}{2} ). Then f_{P} (x) = g(y) for each y ∈ ℝ^{n} such that P − x ∈ T_{P′−y,ϵ}. This is well-defined since there is at most one such y in B(0, \frac{r_{P}}{2} ). Now it is clear that f_{P} is strongly P -equivariant with constant R = ϵ^{−1}. The preservation of smoothness follows directly using the local product structure.

2. If g is strongly P -equivariant and R the constant from Definition 21 then g̃ is transversally locally constant with ϵ = R^{−1}. The preservation of smoothness follows as in 1.

3. The first statement is a direct consequence of 1 and 2 above together with the straightforward observation that f̃_{P} = f and g̃_{P} = g. The second statement follows from the fact that the completion of C_{tlc}^{∞}(Ω_{P} , φ, ℝ) in the Fréchet topology defined by the semi-norms ∥ω∥_{∞} (supremum norm) and ∥ω∥_{i_{1}···i_{k}} = ∥∂_{e_{i_{1}}} ⋯ ∂_{e_{i_{k}}} ω∥_{∞} is C^{∞}(Ω_{P} , φ, ℝ). This is directly seen using the local product structure.

As a consequence of this proposition we have

Theorem 23. f ↦ f_{P} induces graded ring isomorphisms between the transversally locally constant dynamical cohomology H(ℝ^{n}, C_{tlc}^{∞}(Ω_{P} , ℝ)) and the strongly P -equivariant cohomology Ȟ_{P} (ℝ^{n}), and between the dynamical cohomology H(ℝ^{n}, C^{∞}(Ω_{P} , ℝ)) and the weakly P -equivariant cohomology H_{P} (ℝ^{n}).

Remark 24. Let us call a complex continuous function over ℝ^{n} P -equivariant if it lies in the closure of Č_{P} (ℝ^{n}, ℂ) w.r.t the supremum norm. Then f ↦ f_{P} induces an algebra isomorphism between C(Ω_{P}) and this closure. Hence the topological space Ω_{P} is described by the complex continuous P -equivariant functions. The Cech cohomology of Ω_{P} is, however, the cohomology of the sub-complex of strongly P -equivariant forms. The latter was therefore simply called the P -equivariant cohomology (of ℝ^{n}) in [Ke03].

6.1 Multiplicativity of the Ruelle-Sullivan map

The advantage of Theorem 23 is that elements of Ȟ_{P} (ℝ^{n}) have representatives which are forms over ℝ^{n}. These are not only more easily accessible but also allow to use technics from differential topology, as is done in the proposition below. It is used to show that the Ruelle-Sullivan map for ergodic invariant measures of ℝ^{n}-actions on hulls of Delone sets with finite local complexity is multiplicative.

We denote for ω ∈ C^{∞}(X, Λ^{∗})

∥ω∥_{1} = ∫_{X} |ω(x)|dμ(x)

where | · | is a norm on Λ^{∗}.

Proposition 25. Let P be a Delone set with finite local complexity and μ be an ergodic invariant probability measure for the action of ℝ^{n} on Ω_{P}. Let ω be any closed strongly P -equivariant k-form on ℝ^{n}, k ≥ 1. For any ϵ > 0 there exists a closed strongly P -equivariant form ω_{ϵ} which defines the same element as ω in Ȟ_{P} (ℝ^{n}) and such that

∥ω_{ϵ} − τ̃_{φ,μ}(ω)∥_{1} ≤ ϵ.

Here τ̃_{φ,μ} is the expression of the Ruelle-Sullivan map on Ȟ_{P} (ℝ^{n}) and we view τ̃_{φ,μ}(ω) as constant form on ℝ^{n}.

Proof. Let ω be a closed strongly P -equivariant k-form. We let ω_{r} be the form obtained by averaging over r-cubes I_{r} = [−\frac{r}{2}, \frac{r}{2}]^{n}:

ω_{r}(x) = \frac{1}{r^{n}} ∫_{I_{r}} ω(x + y)dy.

By the Birkhoff ergodic theorem ω_{r} converges to τ̃_{φ,μ}(ω) for r → ∞ in the norm ∥ · ∥_{1} [Li01]. Moreover, each ω_{r} is a closed strongly P -equivariant kform. The proposition follows therefore if we can verify that ω_{r} − ω = dG_{r} for a strongly P -equivariant k − 1-form G_{r}.

For any e_{1}, …, e_{k−1} ∈ ℝ^{n} and y ∈ ℝ^{n}, we let g_{y} : ℝ^{n} → Λ^{k−1} be given by

⟨e_{1} ∧ ⋯ ∧ e_{k−1}, g_{y}(x)⟩ = (−1)^{k−1} ∫_{0}^{1} dt⟨e_{1} ∧ ⋯ ∧ e_{k−1} ∧ y, ω(x + ty)⟩.

g_{y} is strongly P -equivariant, as the integral depends only on the values of ω between x and x + y. Furthermore,

⟨e_{0} ∧ ⋯ ∧ e_{k−1}, dg_{y}(x)⟩ = ∑_{i=0}^{k−1}(−1)^{i}e_{i} · ∇⟨e_{0} ∧ ⋯ ê_{i} ⋯ ∧ e_{k−1}, g_{y}(x)⟩

= (−1)^{k−1} ∫_{t=0}^{1} dt (⟨e_{1} ∧ ⋯ ∧ e_{k−1} ∧ y, dω(x + ty)⟩
−(−1)^{k}y · ∇⟨e_{0} ∧ ⋯ ∧ e_{k−1}, ω(x + ty)⟩)

where y · ∇f is the derivative of the function f in the direction y. Since ω is closed the first term vanishes, and using y·∇f(x+ty) = \frac{d}{dt}f(x+ty) the second integral gives ⟨e_{0}∧⋯∧e_{k−1}, ω(x+y)−ω(x)⟩. Hence ω(x+y)−ω(x) = dg_{y}(x) and

G_{r} = \frac{1}{r^{n}} ∫_{I_{r}} g_{y}dy.

G_{r} is thus a strongly P -equivariant form satisfying ω_{r} − ω = dG_{r}.

Theorem 26. Let P be a Delone set with finite local complexity and μ be an ergodic invariant probability measure for the action of ℝ^{n} on Ω_{P}. The Ruelle-Sullivan map

τ_{φ,μ} : Ȟ^{∗}(Ω_{P} , ℤ) → Λℝ^{n}^{∗}

is a homomorphism of graded rings.

Proof. We have seen that the map θ_{φ} is a ring homomorphism. We show that also τ̃_{φ,μ} is multiplicative. Let ω, ρ be two closed strongly P -equivariant forms over ℝ^{n}. Given ϵ > 0 let ω_{ϵ} be as in the statement of Proposition 25. Then

τ̃_{φ,μ}(ωρ) = τ̃_{φ,μ}(ω_{ϵ}ρ)

since (ω − ω_{ϵ})ρ is exact. Furthermore

τ̃_{φ,μ}(ω_{ϵ}ρ) = τ̃_{φ,μ}(τ̃_{φ,μ}(ω_{ϵ})ρ) + ∫_{Ω_{P}} (ω_{ϵ} − τ̃_{φ,μ}(ω_{ϵ}))ρ dμ

and hence

|τ̃_{φ,μ}(ωρ) − τ̃_{φ,μ}(ω)τ̃_{φ,μ}(ρ)| ≤ ∥ω_{ϵ} − τ̃_{φ,μ}(ω)∥_{1}∥ρ∥_{∞} ≤ ϵ∥ρ∥_{∞}

where ∥ρ∥_{∞} = sup_{x} |ρ(x)|. The result follows since ϵ was arbitrary.

References

[AM85] R. Abraham and J.E. Marsden, Foundations of Mechanics, (Addison Wesley, 1985).

[BT82] R. Bott and L.W. Tu, Differential Forms in Algebraic Topology, (Springer-Verlag, 1982).

[Co80] A. Connes, C^{∗}-algèbres et géométrie différentielle, C.R.Acad.Sc. Paris, t. 290, Série A, 599–604.

[C94] A. Connes, Non-Commutative Geometry, (Acad. Press, San Diego, 1994).

[FH99] A.H. Forrest and J. Hunton. The cohomology and K-theory of commuting homeomorphisms of the Cantor set. Ergod. Th. and Dynam. Sys., 19:611–625, 1999.

[FHK02] A.H. Forrest, J. Hunton, and J. Kellendonk, Topological invariants for projection method patterns, Memoirs of the AMS 159(758), x+120 pp (2002).

[Gl76] S. Glasner, Proximal flows, Lecture notes in math, vol 517 Springer, 1976.

[Ke03] J. Kellendonk. Pattern-equivariant functions and cohomology. J. Phys. A 36:1–8 (2003).

[Li01] E. Lindenstrauss, Pointwise theorems for amenable groups, Invent. Math. 146, (2001), 259-295.

[Mo97] R.V. Moody Meyer Sets and their Duals, in The Mathematics of Long-Range Aperiodic Order, pages 403–441, editor R.V. Moody Kluwer Academic Publishers, 1997.

[MS88] C.C. Moore and C. Schochet, Global analysis on foliated spaces, Mathematical Sciences Research Institute Publications, 9, Springer-Verlag, New York, 1988.

[So98] Solomyak, Spectrum of Dynamical Systems Arising from Delone Sets, in Quasicrystals and Discrete Geometry, pages 265–275, editor J. Patera, Fields Institute Monographs, AMS, 1998.